OpenStudy (anonymous):

The pair of points (6, y) and (10, –1) lie on a line with slope 1/4 . What is the value of y? –5 –2 2 5

OpenStudy (anonymous):

slope is given by the formula: m=(y2-y1)/(x2-x1) so, 1/4=(y+1)/(6-10) ==> 1/4=-(y+1)/4 ==> -1=y+1 ==>y=-2